OpenStudy (anonymous):

how to write an equation of the circle with center at the origin and a radius of 11

OpenStudy (anonymous):

The standard form equation for a circle at the origin is x^2+y^2=r^2 where r is the radius. Can you get it now?

OpenStudy (anonymous):

x^2+y^2=11^2 x^2+y^2=121

OpenStudy (anonymous):

R^2=(x-a)^2+(y-b)^2 (x,y) points of the center (a,b) any point on the circumference of the circle this is the general equation just put the given information in it i hope this help

OpenStudy (anonymous):

R is the radius